# Leveraging Climate Services to Build Climate Resilient Power Systems

### 01 May 2026 v1

@@ Abstract
Climate change impacts the energy sector on multiple fronts. Short-term weather variability drives daily and seasonal fluctuations in supply and demand. At the same time, long-term trends and increased frequency of extremes pose risks to infrastructure performance, asset lifetimes, and system adequacy. Representing compound events (e.g., simultaneous low renewable generation and high demand) and spatial correlations across borders is a complex challenge, and uncertainties persist due to uncertainties from different climate models, emission scenarios, and downscaling methodologies.
Traditionally, energy planning has relied on meteorological reanalysis datasets such as ERA5, which offer historical consistency but become less representative as the climate diverges from past conditions. Some operators have begun to use climate projections, but often only a single scenario or model is considered, limiting robust uncertainty assessments. Moreover, many datasets were not public, restricted to national domains, or lacked multivariate coherence and transparency.
The Pan-European Climate Database (PECD4.2), developed in partnership between the European Network of Transmission System Operators for Electricity (ENTSO-E) and the Copernicus Climate Change Service (C3S), marks a step change. It integrates historical reanalysis and six climate models across four shared socio-economic pathways, providing harmonised, openly available datasets tailored for power system studies. PECD4.2 delivers key climate and derived energy variables, at sub-national to national scales and with hourly resolution. Its physical conversion models for wind and solar energy better reflect technological progression than machine learning methods trained on historical data, improving robustness under changing future conditions.
PECD4.2 places special emphasis on multi-variable consistency, bias adjustment for continuity between historical and projected periods, and quantification of uncertainties. This enables more probabilistic and risk-based approaches rather than relying solely on deterministic scenarios. The openly available data and transparent methodologies foster consistency across studies and users.
Despite these advances, challenges remain. Current PECD energy conversion models, particularly for hydropower, require further refinement and country-specific calibration. Progress is slow due to the limited availability and mixed quality of validation datasets. The sharing of detailed, harmonised energy data needs to improve.
Complex processing chains from raw climate data to actionable insights and the lack of standards for integrating climate information into energy planning lengthen lead times for energy-sector adoption. This leads to diverging approaches and variable consideration of climate risks. Closer, more generalised collaboration and communication between climate service providers and energy stakeholders are therefore necessary, as are the development of user-friendly tools for data manipulation and analysis and robust feedback loops.
Standardization and deeper data sharing are essential to advance prospective studies and build climate-resilient, low-carbon power systems capable of withstanding an evolving climate. The openly available PECD, based on state of the art climate models and transparent methodologies, is an essential piece in bridging those gaps.
@@ 1 Introduction
The Earth’s climate is changing primarily due to anthropogenic greenhouse gas (GHG) emissions and this change is fundamentally transforming the boundary conditions within which energy systems are planned and operated [undef]. Globally, the energy sector is responsible for approximately two-thirds of GHG emissions [undefa]. A transition toward low-emissions energy sources is therefore critical and advocates for a high level of electrification in energy production. A comprehensive decarbonization of electricity production is thus at the forefront of climate change mitigation. The Paris Agreement’s goals [undefb] – to limit warming to well below 2°C, preferably 1.5°C – demand that future energy pathways are aligned with scenarios which account for a rapidly changing climate, both in average states and in the increasing frequency and severity of extremes.
In this context, the European Green Deal and subsequent legislation (e.g., ’Fit for 55’ and REPowerEU) stipulate unprecedented efforts towards a net-zero energy system by 2050, to be achieved chiefly through improved energy efficiency and a major increase in the share of renewables [undefc, undefd, undefe]. This legislation aims for a 55% reduction of greenhouse gas emissions in 2030 compared to 1990 for the EU27 and to increase the share of renewable energy in the EU’s energy mix, with a binding target of 42.5% for 2030 and the ambition to reach 45%.
Electricity systems are therefore expected to play a central role in this transition, notably through the electrification of end uses such as transport, heating, and industry. The associated change in electricity demand is expected to be met largely by renewable energy sources, particularly wind, solar photovoltaic, and hydropower. This transformation will make energy systems even more dependent on weather, climate variability, and climate change, introducing new vulnerabilities and requiring new tools to assess future risks related to systems’ operation [undeff, undefg, undefh, undefi, undefj, undefk].
@@ 2 Climate variability and change as key drivers of power systems
Energy systems are sensitive to climate across a wide range of temporal and spatial scales. Short-term weather variability influences daily and seasonal fluctuations in electricity demand and renewable generation, while long-term climate variability and change affect average conditions, extremes, and the statistical distribution of weather events. Energy system assessments require representation and identification of all of these factors, but this is a persistent challenge [undefk].
One of these challenges lies in representing extremes and compound events. Periods of low wind and solar generation coinciding with high electricity demand are of particular concern for system adequacy, as they imply a potential energy shortage. Climate change may alter the frequency, duration, or intensity of such events, but these changes are difficult to assess using short historical records alone. Similarly, compound hazards, such as simultaneous heatwaves and droughts, can affect both electricity demand and hydropower availability [undefl].
Another challenge relates to spatial coherence. Large-scale atmospheric circulation patterns can lead to correlated weather conditions across multiple countries, reducing the effectiveness of geographic diversification of renewable resources [undefm]. Capturing these correlations requires climate datasets with sufficient spatial coverage and physical consistency.
Finally, uncertainty is inherent to climate information, especially when considering future projections. The choice of emission scenarios, climate models, and downscaling methods can substantially influence the projected outcomes. Energy system studies must therefore be designed to account for uncertainties, rather than relying on a single deterministic climate future [undefn].
@@ 3.2 Long-term prospective of system operation
Traditionally, energy system modelling has utilized historical meteorological datasets, such as the ERA5 reanalysis [undefp], to derive input time series for key variables like wind, solar irradiation, and hydropower resources [undefq, undefi, undefj]. While these reanalysis products amalgamate observations into consistent, gridded fields suitable for large-scale modelling, they are by construction representative only of past conditions. Their validity for supporting long-term adequacy studies diminishes as the climate system departs from historical norms [undefn].
ENTSO-E developed the Pan-European Climate Database (PECD), with versions up to 3.0 based solely on reanalysis products [undefp]. Version 3.1 included only limited ad hoc corrections for climate change, by trend-correcting temperature, and did not directly represent future climate scenarios [undefr].
Adequacy studies are particularly demanding in terms of climate data requirements, as they often rely on long time series of weather-driven demand and generation to assess rare but critical events. Climate data used in energy system studies must meet several key requirements (see [undefn] and references therein). First, they should provide long and consistent time series to capture interannual variability and extremes. Second, they must preserve the physical relationships between different climate variables, such as temperature, wind, and precipitation. Third, spatial and temporal resolution should be sufficient to represent regional differences and relevant processes. Finally, transparency and reproducibility are essential, particularly for regulatory and policy-oriented studies.
@@ 3.3 Limitations of current approaches
In the last few years some TSOs adopted climate projection datasets in their studies. These datasets use climate models and GHG emission scenarios to provide potential future climate evolution, most generally up to the end of the century. Examples are the SDDR analysis by RTE mentioned above, their Energy Pathways to 2050 study [undefs], or Elia’s blueprint for the Belgian electricity system [undeft].
@@ 4.1 The new landscape of climate services in Europe
The energy sector has been identified as a key sector for C3S applications since the early stages, leading to two proofs of concept (ECEM [undefu, undefv], CLIM4ENERGY [undefw]), followed by a pre-operational service [undefx].
To respond to the EU Commission and the EU Agency for the Cooperation of Energy Regulators (ACER) ’s request to include climate change in prospective studies, ENTSO-E has partnered with C3S to develop a new, climate-proof version of the PECD [undefn].
@@
Leveraging recent developments in climate services, the PECD has been updated over the last three years to provide a harmonized set of climate-driven energy variables for European transmission system operators (Figure 1) [undefy]. It now includes both historical data based on the ERA5 reanalysis [undefp] and climate projections from 6 CMIP6 global climate models up to 2100, each under 4 Shared Socio-economic Pathways (SSPs) as defined by the IPCC [undef].
@@
It is planned to use this new database in several studies, notably the ENTSO-E ERAA 2026, the TYNDP 2026, and the ongoing RTE’s update of the Energy pathways to 2050 report. Each of these studies has its specific requirements, including differences in the number of climate years that can be processed in their respective power system models (PLEXOS for ENTSO-E, ANTARES simulator for RTE). As a result, climate years subsets selection methodologies have also been proposed [undefz] and will be further developed in the coming months. These aim to identify representative years for a given period, while preserving as much of the variability in the full data ensemble as possible and allowing more explicit alignment with the different study objectives.
@@ 5 Discussion and steps forward
The data processing chains from climate-energy data generation to practical use in energy prospective studies remain complex, which lengthens the time between the production of climate simulations (CMIP, CORDEX), their processing by climate services (C3S), and their adoption and use in the energy sector. As part of the PECD4.2 development, a set of Python notebooks [undefaa] is provided, offering a good starting point for non-specialist users to access, retrieve, and process the data. It will be further developed in the future.
@@ Acknowledgments
The PECD4.2 is available under the as the Copernicus Climate Change Service (2024): Climate and energy related variables from the Pan-European Climate Database derived from reanalysis and climate projections. For more details, download options and the full documentation see [undefy] at https://doi.org/10.24381/cds.f323c5ec

### 04 May 2026 v2

@@ Abstract
Climate change impacts the energy sector on multiple fronts. Short-term weather variability drives daily and seasonal fluctuations in supply and demand. Long-term trends and increased frequency of extremes pose risks to infrastructure performance, asset lifetimes, and system adequacy. Representing compound events and spatial correlations across borders is a complex challenge, and uncertainties persist due to uncertainties from different models, scenarios, and downscaling methodologies.
The Pan-European Climate Database (PECD4.2), developed in partnership between the European Network of Transmission System Operators for Electricity (ENTSO-E) and the Copernicus Climate Change Service (C3S), marks a change in how energy system planning is conducted. The PECD4.2 integrates historical reanalysis and six climate models across four SSP’s, providing harmonised, openly available datasets tailored for power system studies. The physical conversion models for wind and solar energy better reflect technological progression than machine learning methods trained on historical data, improving robustness under changing future conditions.
Despite these advances, challenges remain. Particularly in hydropower modelling and the lack of public harmonised energy datasets that are required to train these models. Complex processing chains from raw climate data to actionable insights and the lack of standardized integration of climate information lengthen lead times for energy-sector adoption. This leads to diverging approaches and variable consideration of climate risks. Closer, more generalised collaboration and communication between climate service providers and energy stakeholders are therefore necessary, as are the development of user-friendly tools for data manipulation and analysis and robust feedback loops.
@@ 1 Introduction
The Earth’s climate is changing primarily due to anthropogenic greenhouse gas (GHG) emissions and this change is fundamentally transforming the boundary conditions within which energy systems are planned and operated [ipcc6wr1]. Globally, the energy sector is responsible for approximately two-thirds of GHG emissions [IEA2021]. A transition toward low-emissions energy sources is therefore critical and advocates for a high level of electrification in energy production. A comprehensive decarbonization of electricity production is thus at the forefront of climate change mitigation. The Paris Agreement’s goals [parisAgreement] – to limit warming to well below 2°C, preferably 1.5°C – demand that future energy pathways are aligned with scenarios which account for a rapidly changing climate, both in average states and in the increasing frequency and severity of extremes.
In this context, the European Green Deal and subsequent legislation (e.g., ’Fit for 55’ and REPowerEU) stipulate unprecedented efforts towards a net-zero energy system by 2050, to be achieved chiefly through improved energy efficiency and a major increase in the share of renewables [greenDeal, SteppingUp, REPowerEU]. This legislation aims for a 55% reduction of greenhouse gas emissions in 2030 compared to 1990 for the EU27 and to increase the share of renewable energy in the EU’s energy mix, with a binding target of 42.5% for 2030 and the ambition to reach 45%.
Electricity systems are therefore expected to play a central role in this transition, notably through the electrification of end uses such as transport, heating, and industry. The associated change in electricity demand is expected to be met largely by renewable energy sources, particularly wind, solar photovoltaic, and hydropower. This transformation will make energy systems even more dependent on weather, climate variability, and climate change, introducing new vulnerabilities and requiring new tools to assess future risks related to systems’ operation [Jerez2015, Tobin2016, Tobin2018, Dubus2018, Bloomfield2021, Craig2022].
@@ 2 Climate variability and change as key drivers of power systems
Energy systems are sensitive to climate across a wide range of temporal and spatial scales. Short-term weather variability influences daily and seasonal fluctuations in electricity demand and renewable generation, while long-term climate variability and change affect average conditions, extremes, and the statistical distribution of weather events. Energy system assessments require representation and identification of all of these factors, but this is a persistent challenge [Craig2022].
One of these challenges lies in representing extremes and compound events. Periods of low wind and solar generation coinciding with high electricity demand are of particular concern for system adequacy, as they imply a potential energy shortage. Climate change may alter the frequency, duration, or intensity of such events, but these changes are difficult to assess using short historical records alone. Similarly, compound hazards, such as simultaneous heatwaves and droughts, can affect both electricity demand and hydropower availability [Biewald2025].
Another challenge relates to spatial coherence. Large-scale atmospheric circulation patterns can lead to correlated weather conditions across multiple countries, reducing the effectiveness of geographic diversification of renewable resources [vanDuinen2025]. Capturing these correlations requires climate datasets with sufficient spatial coverage and physical consistency.
Finally, uncertainty is inherent to climate information, especially when considering future projections. The choice of emission scenarios, climate models, and downscaling methods can substantially influence the projected outcomes. Energy system studies must therefore be designed to account for uncertainties, rather than relying on a single deterministic climate future [Dubus2022].
@@ 3.2 Long-term prospective of system operation
Traditionally, energy system modelling has utilized historical meteorological datasets, such as the ERA5 reanalysis [hersbach2020era5], to derive input time series for key variables like wind, solar irradiation, and hydropower resources [Craig2018, Dubus2018, Bloomfield2021]. While these reanalysis products amalgamate observations into consistent, gridded fields suitable for large-scale modelling, they are by construction representative only of past conditions. Their validity for supporting long-term adequacy studies diminishes as the climate system departs from historical norms [Dubus2022].
ENTSO-E developed the Pan-European Climate Database (PECD), with versions up to 3.0 based solely on reanalysis products [hersbach2020era5]. Version 3.1 included only limited ad hoc corrections for climate change, by trend-correcting temperature, and did not directly represent future climate scenarios [PECDupdate].
Adequacy studies are particularly demanding in terms of climate data requirements, as they often rely on long time series of weather-driven demand and generation to assess rare but critical events. Climate data used in energy system studies must meet several key requirements (see [Dubus2022] and references therein). First, they should provide long and consistent time series to capture interannual variability and extremes. Second, they must preserve the physical relationships between different climate variables, such as temperature, wind, and precipitation. Third, spatial and temporal resolution should be sufficient to represent regional differences and relevant processes. Finally, transparency and reproducibility are essential, particularly for regulatory and policy-oriented studies.
@@ 3.3 Limitations of current approaches
In the last few years some TSOs adopted climate projection datasets in their studies. These datasets use climate models and GHG emission scenarios to provide potential future climate evolution, most generally up to the end of the century. Examples are the SDDR analysis by RTE mentioned above, their Energy Pathways to 2050 study [EnergyPathways2050], or Elia’s blueprint for the Belgian electricity system [BelgianBlueprint].
@@ 4.1 The new landscape of climate services in Europe
The energy sector has been identified as a key sector for C3S applications since the early stages, leading to two proofs of concept (ECEM [Troccoli2018, Goodess2019], CLIM4ENERGY [Bartk2019]), followed by a pre-operational service [Dubus2023].
To respond to the EU Commission and the EU Agency for the Cooperation of Energy Regulators (ACER) ’s request to include climate change in prospective studies, ENTSO-E has partnered with C3S to develop a new, climate-proof version of the PECD [Dubus2022].
@@
Leveraging recent developments in climate services, the PECD has been updated over the last three years to provide a harmonized set of climate-driven energy variables for European transmission system operators (Figure 1) [PECD42]. It now includes both historical data based on the ERA5 reanalysis [hersbach2020era5] and climate projections from 6 CMIP6 global climate models up to 2100, each under 4 Shared Socio-economic Pathways (SSPs) as defined by the IPCC [ipcc6wr1].
@@
It is planned to use this new database in several studies, notably the ENTSO-E ERAA 2026, the TYNDP 2026, and the ongoing RTE’s update of the Energy pathways to 2050 report. Each of these studies has its specific requirements, including differences in the number of climate years that can be processed in their respective power system models (PLEXOS for ENTSO-E, ANTARES simulator for RTE). As a result, climate years subsets selection methodologies have also been proposed [vanDuinen2026] and will be further developed in the coming months. These aim to identify representative years for a given period, while preserving as much of the variability in the full data ensemble as possible and allowing more explicit alignment with the different study objectives.
@@ 5 Discussion and steps forward
The data processing chains from climate-energy data generation to practical use in energy prospective studies remain complex, which lengthens the time between the production of climate simulations (CMIP, CORDEX), their processing by climate services (C3S), and their adoption and use in the energy sector. As part of the PECD4.2 development, a set of Python notebooks [PECDtraining] is provided, offering a good starting point for non-specialist users to access, retrieve, and process the data. It will be further developed in the future.
@@ Acknowledgments
The PECD4.2 is available under the as the Copernicus Climate Change Service (2024): Climate and energy related variables from the Pan-European Climate Database derived from reanalysis and climate projections. For more details, download options and the full documentation see [PECD42] at https://doi.org/10.24381/cds.f323c5ec